1)Find all real numbers that satisfy the equation: cos x = square root 3 / 2.
2)Solve the equation: 6 cos(Alpha)+3=0 for 0degree 3)Find all real numbers in the interval[0,2pi) that satisfy the equation: cos^2x+2cosx+1=0.
4)Find all values of x in the interval [0degree,360 degree)that satisfy the equation. Round approximate answer to the nearest tenth of a degree: sin^2x+8sinx-4=0.
5)Find all angles in degrees that satisfy the equation: cos alpha=-0.609

Found 2 solutions by lynnlo, ikleyn:
Answer by lynnlo(4176)
Answer by ikleyn(53427)

.
1) Find all real numbers that satisfy the equation: cos x = square root 3 / 2.
2) Solve the equation: 6 cos(Alpha)+3=0 for 0 degree < alpha < 360 degree.
3) Find all real numbers in the interval[0,2pi) that satisfy the equation: cos^2x+2cosx+1=0.
4) Find all values of x in the interval [0degree,360 degree)that satisfy the equation.
Round approximate answer to the nearest tenth of a degree: sin^2x+8sinx-4=0.
5) Find all angles in degrees that satisfy the equation: cos alpha=-0.609
~~~~~~~~~~~~~~~~~~~~~~~~~~~

NEVER pack more than one problem per post.